calculate gtt/min using macrotubing (20gtts/mL) to infuse 150 mL/hr, then 100mL/hr
Answer #1

Order dose to infuse 150 mL/hr.

Drop factor 20gtts/mL

Drop calculation formula - (Total fluid volume in ml \times Drop factor) \div Time in minute

= (150 ml \times 20gtt/ml) \div 60 min

= 50 gtt/min

Answer - The flow rate will be 50 gtt/min

If Order dose to infuse 100 mL/hr.

Then, = (100 ml \times 20gtt/ml) \div 60 min

= 33 gtt/min

Answer - The flow rate will be 33 gtt/min